OpenStudy (anonymous):

Tracy ran for seven-twelfths hours and Reina ran for one-third hours. How many more hours did Tracy run? one-fourth 1 one-third one-sixth

OpenStudy (anonymous):

the question is asking how much more is 7/12 than 1/3 ? get them both with a common denominator (they need the same number on the bottom) 1/3 = 4/12 so we can see that 7/12 is 3/12 more than 4/12 because 7/12 - 4/12 = 3/12 3/12 is the same as 1/4 comment if you dont understand how i changed the bottom numbers to 12 and back

OpenStudy (anonymous):

so its will be D

OpenStudy (anonymous):

no A, 1 fourth is 1/4